Review Conceptual Example 3 before starting this problem. The amplitude of a transverse wave on a string is 3.63 cm. The ratio of the maximum particle speed to the speed of the wave is 3.90. What is the wavelength (in cm) of the wave?
Answer #1

The relation between velocity,frequency and wavelength is

V = f*wavelength

V / wavelength = (A*2*pi*f)/(f*I)

3.9 = (A*2*pi)/ I

I = (2*3.63*3.14)/ 3.9

= 5.84 cm